OpenStudy (peterg2001):

Which of the following is an example of the commutative property? (a + b) + c = (b + a) + c (a + b) + c = a + (b + c) a(b + c) = ab + ac

OpenStudy (peterg2001):

@Monkey4224

OpenStudy (gabylovesyou):

An example of commutative property is "3 + 4 = 4 + 3". Sooooooooo.... what do u think the answer is ?

OpenStudy (anonymous):

(a + b) + c = (b + a) + c

OpenStudy (gabylovesyou):

Commutative property just means that both side should equal the same no matter the order. lets say a = 1 , b = 2 and c = 3. Try A: (1 + 2) + 3 = (2 + 1) + 3 both sides equal 6 soon A is correct :D

OpenStudy (gabylovesyou):

sooo*